# A Two-Stage Fourth-Order Implicit Scheme for Stiff problems

### v1

@@ Abstract
This paper presents a novel two-stage fourth-order implicit scheme designed to overcome the limitations of existing spatiotemporal coupling two-stage fourth-order methods, which have thus far been confined to explicit frameworks. In such frameworks, computational efficiency is severely constrained by the CFL condition when addressing stiff problems, and they fundamentally conflicts with spatiotemporal coupled methods whose core advantage lies in embedding stiff source terms. By fully leveraging temporal derivatives of physical variables within a rigorously derived compact mathematical framework, the scheme achieves fourth-order temporal accuracy in only two stages. Furthermore, a sufficient condition for A-stability is established through systematic theoretical and numerical investigation, and a Newton iterative procedure is provided to accelerate convergence. Extensive numerical experiments on classical stiff problems confirm the method’s effectiveness and competitiveness.
Key words: two-stage fourth-order; temporal derivatives; implicit scheme; A-stability; stiff problems; Newton iteration
@@ 1 Introduction
The generalized Riemann problem (GRP) fundamentally differs from the classical Riemann problem through its use of spatially varying initial data, which generates non-self-similar solutions with non-zero time derivatives. This critical advancement enables the natural incorporation of physical source terms and other essential effects (e.g., continuous entropy variation) into the temporal evolution, establishing GRP as an indispensable framework for capturing complex physical mechanisms. The simplest GRP form replaces piecewise constant with piecewise linear initial data in homogeneous systems, first explored by Kolgan [1] and further developed by van Leer [2, 3]. Toro and Titarev [4] later extended GRP to include source terms using the Lax-Wendroff procedure, forming the basis of arbitrarily high-order (ADER) schemes [5, 6] applicable to hydrodynamics, magnetohydrodynamics [7], reaction-diffusion [8], and stiff systems [9], though these linearized solvers struggle under extreme thermodynamic conditions. Ben-Artzi et al. [10, 11, 12] were the first to develop the GRP solver for compressible fluids, which can be considered as the nonlinear version of the ADER solver. By incorporating thermodynamic effects, the GRP solver can accurately resolve strong waves, making it valuable under extreme conditions [13]. The GRP solver has since been successfully applied to duct flow [14], reactive flow [15], shallow water equations [16, 17], multi-material and multiphase flow [18, 19, 20, 21], magnetohydrodynamics [22], relativistic hydrodynamics [23, 24, 25], blood flow modeling [26], and radiation hydrodynamics [27].
The inherent spatiotemporal coupling in fluid dynamics necessitates numerical algorithms that preserve such correlations, which conventional Riemann solvers fail to achieve as they only provide piecewise constant states for constructing instantaneous Cauchy fluxes. This limitation leads to error accumulation over time in flows with strong discontinuities. In contrast, generalized Riemann solvers (e.g., ADER or GRP) provide temporal derivatives of physical quantities, enabling construction of time-interval fluxes with inherent spatiotemporal coupling. These fluxes exhibit Lipschitz continuity with respect to boundary perturbations, allowing temporal effects to enhance spatial regularity. Consequently, errors from strong discontinuities can be directly controlled through the time step, enabling accurate long-term resolution while establishing consistency between mathematical weak solutions and physical integral balance laws [28, 29].
Runge-Kutta methods, originally developed by Carl Runge [30] in 1895 and Martin Wilhelm Kutta [31] in 1901, constitute a family of classical numerical techniques for solving initial-value problems of ordinary differential equations that comprise both explicit and implicit forms. The explicit variants, exemplified by the classical fourth-order RK4 scheme, achieve high-order accuracy through weighted averages of function values at multiple stages, offering implementation simplicity without requiring calculation of higher-order derivatives. However, their stability remains constrained by the CFL condition, rendering them inefficient for stiff problems. In contrast, Implicit Runge-Kutta methods proposed by Kuntzmann [32] and Butcher [33], such as Radau and Gauss-type schemes, compute slopes at future time steps by solving systems of nonlinear equations. Although computationally more expensive, they exhibit superior numerical stability, making them particularly suitable for stiff differential equations.
@@
Recently, Li et al. developed a two-stage fourth-order time advancement method [34], which makes full use of the temporal derivatives of physical quantities, enabling the scheme to achieve fourth-order time accuracy in only two stages. This method successfully combines the simplicity advantages of Runge-Kutta-type methods with the spatiotemporally coupling properties of the GRP solver. Then they details the characteristics of this class of methods [35]: (i) Computational efficiency: Under the same conditions, its computational efficiency is twice that of the Runge-Kutta method, which requires four stages to achieve fourth-order time accuracy [36]; (ii) Compactness: Since the number of time advancement steps is reduced by half, the stencil size is also halved; the spatial-temporal coupled HWENO-type reconstruction [37, 38] can make the computational scheme more compact; (iii) Stability: It has been proven that its stability region is larger than that of Runge-Kutta methods [39]; (iv) Compatibility: The method can be easily combined with other methods, such as the GKS solver [40], the multi-moment method [41], and the CRP reconstruction technique [42].
However, despite the aforementioned advantages, current spatiotemporally coupled two-step fourth-order methods still suffer from significant limitations. On the one hand, such methods are currently confined to explicit frameworks. As a result, their time step is severely constrained by the CFL stability condition, and this restriction in turn drastically reduces their computational efficiency when dealing with stiff problems. On the another hand, the stiffness of a problem is often predominantly manifested in its source terms. For instance, in problems involving multi-scale physics like combustion or chemically reacting flows, the source terms can introduce significant stiffness. The key advantage of the GRP solver over the conventional Riemann solver lies in its ability to provide the temporal derivatives of physical quantities, and this derivative information inherently encodes the effects of the source terms from the governing equations. However, a critical paradox emerges: this same temporal derivative data is utilized in two distinct ways: it is essential for constructing the time-interval fluxes and serves as the cornerstone for building the two-stage fourth-order explicit scheme. The former (time-interval fluxes) can successfully incorporate information from stiff source terms, whereas the latter (the explicit scheme itself) is fundamentally unsuitable for solving stiff problems due to its explicit nature. This inherent incompatibility points to a significant limitation in the current methodological framework. Inaddition, the incorporation of temporal derivatives may allows the proposed method to achieve not only higher computational accuracy but also a more compact formulation compared to existing implicit schemes. Based on the above considerations, we conclude that there is an urgent need to develop two-stage fourth-order implicit methods to address multi-scale problems such as those involving strong stiffness.
As the foundational component of this research on spatiotemporally coupled two-stage fourth-order implicit schemes, hereafter referred to as Implicit TSFO, we develop an implicit framework that attains fourth-order accuracy in two stages through the incorporation of temporal derivatives of physical quantities. Section 2 presents the detailed formulation of the scheme via rigorous mathematical derivation. This is followed in Section 3 by a thorough analysis of the stability conditions of the implicit method. Section 4 outlines the implementation procedure based on Newton iteration. In Section 5, the efficacy of the proposed scheme is demonstrated through several canonical numerical examples, supplemented by comparative studies with classical fourth-order implicit and explicit Runge-Kutta methods. Finally, Section 6 concludes the paper and suggests potential directions for future research.
2 The Construction of the Implicit TSFO Scheme
@@
where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist. We write (2.1) as
@@
where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$, denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
@@
We construct a two-stage fourth-order implicit scheme by utilizing the values of physical quantities along with their temporal derivatives. The proposed scheme is formulated as follows:
Stage 1. We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$ and $t_{n+\frac{1}{2}}$ to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
@@
where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy in Stage 1.
Stage 2. We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$ to obtain the solution at the next time level ${\bf u}^{n+1}$,
@@
where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy in $\bf Stage2$.
@@
$\quad\quad\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
@@
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+\mathcal{O}[(\Delta t)^{5}].$
Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following equation relationship,
$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384},\end{cases}$ (2.17)
Solve the above system of equations, we obtain the unique solution
@@
$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24},\end{cases}$ (2.20)
@@
$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C.\end{cases}$ (2.21)
where $C$, $D$ are real parameters to be determined. Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired two-stage fourth-order implicit (e.g., Implicit TSFO) scheme as follows:
@@
$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
3 The Stability Analysis of the Implicit TSFO Scheme

### v2

# An Implicit Two-Stage Fourth-Order Temporal Discretization Scheme of Lax-Wendroff Type for Hydrodynamic Problems with Stiff Source Terms Part I: Formulation, Stability Analysis, and Newton’s Iteration

@@ Abstract
Existing two-stage fourth-order (TSFO) methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] are confined to explicit frameworks and can only be applied to homogeneous hydrodynamic problems describing idealized states. More importantly, the ability of these methods to achieve fourth-order time accuracy in merely two stages relies on utilizing temporal derivatives of physical quantities, which are obtained through generalized Riemann problem (GRP) solvers. While the advantage of the GRP solvers [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] over traditional Riemann solvers [28, 29] lies precisely in their ability to naturally incorporate the effects of source terms. Specifically, by employing the Lax-Wendroff methodology, the temporal derivatives of physical quantities are converted into spatial derivatives of fluxes and source terms through the governing equations, thereby avoiding errors arising from operator splitting. Therefore, to extend the applicability of TSFO methods to hydrodynamic problems with stiff source terms that describe real physical phenomena, and to fully leverage the advantages of these GRP solvers [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], there is an urgent need to extending TSFO methods from explicit frameworks to implicit ones. The present work lays a solid theoretical foundation and provides a concrete implementation pathway for this research direction. Firstly, this paper derives, through rigorous mathematical analysis, an implicit temporal discretization scheme that achieves fourth-order accuracy in just two stages. Subsequently, sufficient conditions for A‑stability of the proposed implicit scheme are established via systematic theoretical analysis, and a corresponding Newton iteration procedure is formulated to accelerate convergence. Finally, the proposed implicit scheme is extensively validated using classical stiff benchmark problems, confirming its effectiveness and competitiveness. Numerical results demonstrate that the proposed scheme indeed achieves fourth-order time accuracy in two stages and, compared to the classical fourth‑order implicit Runge‑Kutta method, offers a larger stable time step and smaller convergence error.
Key words: two-stage fourth-order (TSFO); generalized Riemann problem (GRP); Lax-Wendroff methodology; implicit scheme; stiff source term; A-stability
@@ 1 Introduction
Hydrodynamic problems with stiff source terms establish a fundamental mathematical framework for describing complex physical phenomena in the real world. This framework distinguishes physical processes according to their characteristic time scales: the hydrodynamic equations govern slow, wave-dominated macroscopic dynamics, while stiff source terms capture fast, localized microscopic interactions. Classical examples include shock-reaction coupling in detonation waves, pollutant-photochemical reactions in atmospheric transport, shock-radiation coupling in astrophysical flows, and the co-evolution of magnetohydrodynamic waves with atomic-scale processes in fusion plasmas. By contrast, homogeneous hydrodynamic equations reduce such complex flow systems in real physical environments to purely convective systems characterized by a single time scale. In numerical implementation, the latter, which involves only one time scale, can typically be solved efficiently using explicit schemes. The former, however, contains multiple distinct fast and slow time scales within the same system. If a purely explicit scheme is applied, the time step constrained by the CFL stability condition will be limited by the smallest time scale, leading to a sharp decline in computational efficiency. Therefore, such systems often require implicit or semi-implicit schemes, in which slowly varying fluid convection is treated explicitly while rapidly varying stiff source terms are handled implicitly, enabling efficient and stable numerical simulation across multiple scales.
The Lax-Wendroff method, serving as a discrete implementation of the Cauchy-Kovalevskaya theorem, establishes and popularizes the fundamental paradigm of directly coupling temporal and spatial derivatives through the governing equations to achieve high-order discretization. It inherently supports the embedding of stiff source terms. Its core concept can be metaphorically described as predicting temporal evolution from spatial structures. Specifically, the method employs a Taylor expansion in time and, by means of the governing equations, transforms higher-order temporal derivatives into combinations of spatial derivatives that incorporate both flux terms and source terms. As a result, temporal evolution information from source terms is automatically integrated into the construction of local time-advancement operators, thereby enabling a unified description of multiscale physical processes and avoiding numerical errors introduced by operator-splitting methods that separately treat fluxes and sources. This characteristic makes the Lax-Wendroff framework particularly suitable for developing implicit high-order schemes, offering an effective mathematical framework for solving complex multiscale flow problems in real-world physics.
The philosophical principle of Lax-Wendroff, which couples space and time through the governing equations themselves, has become a cornerstone for developing advanced numerical methods. One of its key applications is the construction of high‑accuracy generalized Riemann problem (GRP) solvers. For instance, Men’shov’s modified GRP solver [14] enhances the resolution of complex wave interactions; ADER solver [15, 16, 17, 18, 19, 20, 21] enables arbitrary‑order accuracy in both space and time; and Ben-Artiz’s nonlinear GRP solver [22, 23, 24, 25, 26, 27] provides precise treatment of strong discontinuities in compressible flows. These modern descendants are capable of handling hydrodynamic systems with stiff source terms within a unified, unsplit framework, thereby avoiding operator‑splitting errors and preserving the asymptotic properties of complex multiscale physical systems. Furthermore, such GRP solvers are essential for simulating real‑world flow fields containing shocks, turbulence, material interfaces, and intricate physical phenomena, where the regularity of the flow is typically very weak. This stems from the following reasons. Classical Riemann solvers like HLL solver [28] and Roe solver [29] are limited to resolving instantaneous states to construct Cauchy-type instantaneous fluxes, often leading to accumulated numerical errors near strong discontinuities. In contrast, GRP solvers incorporate temporal derivative information to capture local evolutionary trends, enabling construction of temporal-interval fluxes. According to finite-volume principles, these temporal-interval fluxes compensate for spatial regularity deficiencies through their built-in temporal evolution, allowing accurate resolution of strong discontinuities [30, 31].
@@
The philosophy of the Lax-Wendroff method is applicable not only to the development of high‑accuracy GRP solvers, but also to the construction of high‑order numerical schemes. The classical Lax‑Wendroff procedure can achieve arbitrarily high‑order temporal accuracy in a single step, but it requires repeated differentiation of the governing equations. In contrast, Runge‑Kutta methods [32, 33, 34, 35] achieve high‑order temporal accuracy without differentiating the governing equations, yet they involve multi‑stage time integration, which inevitably enlarges the computational stencil. To effectively balance these two strategies, Li et al. [1] proposed a two‑stage fourth‑order (TSFO) temporal discretization scheme that makes full use of the temporal derivative information of physical quantities and attains fourth‑order accuracy in only two stages. The research team later systematically elaborated the advantages of such kind of methods [3]: (1) Computational efficiency: Under comparable conditions, its computational efficiency is twice that of a four‑stage, fourth-order Runge‑Kutta method [4]; (2) Compactness: Because the number of time‑advancement stages is halved, the size of the computational stencil is correspondingly reduced; GRP-solver-based spatiotemporally coupled HWENO‑type reconstruction techniques [5, 6] further enhance the compactness of the scheme; (3) Stability: Its stability region has been proven to be larger than that of Runge‑Kutta methods [7]; (4) Compatibility: The method can be flexibly combined with various other numerical techniques, such as the GKS solver [8], the multi‑moment method [10], and the CRP reconstruction technique [11].
Nevertheless, despite these advantages, existing TSFO methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] remain confined to explicit frameworks. As a result, they are only applicable to homogeneous hydrodynamic systems and cannot effectively address problems involving stiff source terms. What is more important, the capability of such schemes to achieve fourth-order temporal accuracy in merely two stages originates from their design, which fully exploits temporal derivative information of the physical quantities, which are derived from Lax-Wendroff-based GRP solvers. And the key advantage of these solvers lies precisely in their ability to integrate spatial derivatives of fluxes and source terms simultaneously into the time derivatives of physical quantities through the governing equations, all within a unified framework, thereby eliminating numerical errors that typically arise from operator splitting. Consequently, there is a clear and pressing need to develop an implicit TSFO temporal discretization scheme. The development of such a scheme would not only extend the applicability of TSFO methods from homogeneous hydrodynamic problems to systems with stiff source terms, but would also fully harness the Lax-Wendroff methodology’s inherent capacity for seamless stiff-source integration.
As the theoretical foundation and innovative core of this study, the paper is structured as follows. Chapter 2 introduces a rigorously derived implicit temporal discretization scheme that achieves fourth-order accuracy in only two stages by fully utilizing the temporal derivative information of the physical quantities. Chapter 3 establishes sufficient conditions for the A-stability of the proposed scheme through systematic theoretical analysis. Chapter 4 further develops a Newton‑iteration formulation of the scheme to accelerate convergence. Chapter 5 validates the accuracy and stability of the method using a series of classical stiff benchmark problems, and confirms its effectiveness and competitiveness by comparing its performance with classical implicit and explicit Runge‑Kutta schemes. Finally, Chapter 6 summarizes the work and outlines potential directions for future research.
2 Formulation for the Implicit TSFO Scheme
@@
where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist.
In the follows, a two-stage fourth-order implicit scheme for system (2.1) is derived through rigorous mathematical analysis. (2.1) is rewritten as
@@
where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$. Denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
@@
The proposed two-stage fourth-order implicit scheme is given as follows:
Stage 1. The solution at the intermediate time level, $\mathbf{u}^{n+\frac{1}{2}}$, is obtained by utilizing the physical quantities and their temporal derivatives at the time levels $t_{n}$ and $t_{n+\frac{1}{2}}$,
@@
where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy.
Stage 2. The solution at the next time level, $\mathbf{u}^{n+1}$, is obtained by utilizing the physical quantities and their temporal derivatives at the time levels $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$,
@@
where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy.
@@
$\quad\quad\quad\quad\quad\;\;\;+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
@@
$\quad\quad\quad\quad\;\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\;\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\;\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\;\;\;+\mathcal{O}[(\Delta t)^{5}].$
Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following system of equations,
$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384}.\end{cases}$ (2.17)
Solve the above system of equations, we obtain the unique solution,
@@
$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24}.\end{cases}$ (2.20)
@@
$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C,\end{cases}$ (2.21)
where $C$, $D$ are real parameters to be determined. Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired implicit two-stage fourth-order (Implicit TSFO) temporal discretization scheme as follows:
@@
$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where $C$, $D$ are real parameters to be determined.
3 Stability Analysis for the Implicit TSFO Scheme